5.
a) How many six-letter “words” can be formed from the letters in the word
EUCLID if each letter is used once?
b) How many six-letter “words” can be formed from the letters in the word EUCLID if
each letter is used once and the “word” must begin with a consonant?
c) If the six letters are arranged randomly, what is the probability that the resulting
“word” will begin with a consonant?

Answer by ewatrrr(24785)



Hi
EUCLID (6 letters, two vowels, no repeats)
a) 6! = 6*5*4*3*2*1 = 720 ways
b) 2*5*4*3*2*1 = 240 ways
c) P= (2*5*4*3*2*1)/(6*5*4*3*2*1) =  240/720 = 2/6 0r .3333

Wish You the Best in your Studies.